An elevator ascends from the ground with uniform speed. At time T1 a boy drops a marble through a hole in the floor. The marble hits the ground T2 seconds later. Find the height of the elevator at time T1.(T1=T2=4s)

Grade:12th pass

2 Answers

downey zer
6 years ago
Let the speed of the elevator be v
Then initially marble was also travelling at speed v.
when the marble fell from the hole the elevator was at height v*T1 as it was travelling at speed v for T1 time
now use 
h=ut+0.5at^2
where h = – v*T1
a= – g, u=v
t=T2
putting T1=T2=4
u will get v=10
so height =v*T1=40
downey zer
6 years ago
please approve the answer if it helped